the sum of the reciprocal of a number and the reciprocal of 3 less than the number is 8 times the reciprocal of the original number. find the original number.
choices (in case this helps)
a) 7/2
b) -2
c) 8/5
d) 21/8

Answer by mananth(16949)

let number be x
reciprocal = 1/x
reciprocal of 3
=1/3
(1/3)-x
(1/x) + (1/(x-3) = (8/x)
multiply equation by x(x-3)
(x-3) +x=8(x-3)
x-3+x=8x-24
6x=21
x=7/2